# In Context Learning and Reasoning for Symbolic Regression with Large Language Models

### v2

@@ 4 Discussion and Conclusions
SR programs that optimize for speed aim to generate expressions quickly. In contrast, our proposed method emphasizes informed optimization, leveraging contextual information more effectively. The clearest cases of “leveraging the context” occurred when the first guess included the target model among three expressions. But even when the search took longer, we found incorporation of the context, data, and scratchpad to be helpful for improving the quality of generated expressions. However, including noisy data in the context sometimes undermined the search, as did including lower-quality scientific context. Nonetheless, this comes with great computational expense - especially since large datasets and long reasoning chains require so many tokens. GPT-4o is an upgraded version of GPT-4 that offers improved performance, being both faster and more cost-effective [93]. Consistent with observations made by others who have tested it on various tasks [94, 95], we also noted improvements in the quality of generated expressions and the effectiveness of scratchpad analysis through reasoning for symbolic regression.
@@
In general, we found natural language to be a rather clumsy interface for controlling expression length. Different prompts and feedback mechanisms led to distributions of expressions with varying length. Classical approaches that incorporate expression length into measures of fitness or score are certainly more precise for controlling length, even if expression length is an imperfect measure of parsimony and meaningfulness in SR [96]. We ran separate, focused tests to evaluate the effect of prompting on expression length with the GPT-4 model. However, it did not obey instructions that requested, for example, “expressions of length 17.” Even with a scratchpad available, it failed to both measure complexity of an expression accurately and to generate expressions of the target length. The LLM is better-suited for creative generation, while deterministic Python tools are more effective (and cheaper) at procedural tasks such as counting operators and variables to calculate complexity.
Testing our approach on equation rediscovery using language models invariably involves a form of “test set leakage” since these expressions are on Wikipedia and countless additional Internet sources. Indeed, the data are publicly available, as well (appearing, for example, as tables in the .pdf documents of the historical papers), though we think it unlikely that LLMs trained on natural language would devote a sufficient fraction of their network to memorize these datasets. We found strong evidence of this when the scratchpad revealed the LLM naming Kepler’s law before seeing the data. Nonetheless, we found evaluating the model outputs to be informative.
@@
Moreover, we foresee a use case for scientists trying to solve a mystery about their data while having a great deal of context to potentially include. Our experiments included $<$100 words of context, but this context could include experimental details, instrument specifications, and related literature, pushing the limits of the LLM’s context window, or going beyond the context window in a retrieval-augmented generation scheme [97].
@@
The recent study by Shojaee and coworkers [37] demonstrated strong performance in generating scientific expressions using LLMs, using an approach that generated more structured outputs than our approach. Rather than rediscovering scientific laws as standalone expressions, they tasked the LLMs to “autocomplete” a structured Python function within the context of an optimization problem. They also designed new benchmark scenarios and datasets to highlight the strengths of this approach. This approach effectively leverages the strengths of language models in reasoning about well-defined problem statements. However, the extent to which its success can be attributed to prior exposure to similar questions during training remains unclear without a scratchpad to gain insights into its reasoning steps. A true blind test with these language models would be to rediscover an unpublished or recently-published scientific law, using an LLM with a knowledge cutoff predating the seminal publication.
There are two ways to implement in-context learning in LLMs through prompts. One is the Few-Shot prompting method, where we condition the model on a few specific examples related to the task that helps the model understand and perform the task more accurately. The other is the Zero-Shot prompting method, where the output relies solely on template-based prompts without specific task examples, allowing the model to infer how to handle the task from general instructions. Our approach does not provide examples of symbolic regression procedures within the prompts. Instead, we guide the model to engage in freeform chain-of-thought reasoning about the context as it prepares suggestions for new equations.
@@
Biasing the search space in a standard SR program can be challenging and requires advanced software and coding skills. Interdisciplinary work demands significant time and resources from researchers. Natural language interfaces in LLMs can help reduce some of these barriers by making program execution more accessible, even without expertise in software development. Well-crafted prompts empower language models to perform diverse tasks, allowing them to adapt to different contexts and objectives. We can guide these models through prompts to generate responses that align with our specific needs, whether solving complex problems, developing creative content, or analyzing data.
@@ 5 Supporting Information
Figure 19: GPT-4-generated model does not follow theory on extrapolation. It gives a negative value for adsorption loading at lower pressures.

### v3

@@ 4 Discussion and Conclusions
SR programs that optimize for speed aim to generate expressions quickly. In contrast, our proposed method emphasizes informed optimization, leveraging contextual information more effectively. The clearest cases of “leveraging the context” occurred when the first guess included the target model among three expressions. But even when the search took longer, we found that incorporating the context, data, and scratchpad helped improve the quality of generated expressions. However, including noisy data in the context sometimes undermined the search, as did including lower-quality scientific context. Nonetheless, this comes with great computational expense, especially since large datasets and long reasoning chains require so many tokens. GPT-4o is an upgraded version of GPT-4 that offers improved performance, being both faster and more cost-effective [48]. Consistent with observations made by others who have tested it on various tasks [76, 102], we also noted improvements in the quality of generated expressions and the effectiveness of scratchpad analysis through reasoning for symbolic regression.
@@
In general, we found natural language to be a somewhat clumsy interface for controlling expression length. Different prompts and feedback mechanisms led to distributions of expressions with varying length. Classical approaches that incorporate expression length into measures of fitness or score are certainly more precise for controlling length, even if expression length is an imperfect measure of parsimony and meaningfulness in SR [28]. We ran separate, focused tests to evaluate the effect of prompting on expression length with the GPT-4 model. However, it did not obey instructions that requested, for example, “expressions of length 17.” Even with a scratchpad available, it failed to both measure the complexity of an expression accurately and to generate expressions of the target length. The LLM is better-suited for creative generation, while deterministic Python tools are more effective (and cheaper) at procedural tasks such as counting operators and variables to calculate complexity.
Testing our approach on equation rediscovery using language models invariably involves a form of “test set leakage” since these expressions are on Wikipedia and countless additional Internet sources. While the data are publicly accessible, for instance, in tables within historical research papers in PDF format, we think it is unlikely that LLMs trained primarily on natural language would devote a significant portion of their network to memorize these specific datasets. We found strong evidence when our prompts triggered LLM’s memorization of Kepler’s Law, where scratchpad revealed that both GPT-4 and GPT-4o models associate the variable names and data with Kepler’s Law, and it not only guesses the right answer in the first iteration, it names Kepler’s Law in its justification. Perhaps because this relationship is routinely taught in high school and college physics courses, and also the tabulated data is in Wikipedia, and thus likely to be more represented in OpenAI’s training data than others. In our analysis of adsorption datasets, we did not find such instances. The model proposed commonly referenced adsorption isotherms from the literature and recommended them based on patterns in the dataset; however, it never linked the raw data numbers to an adsorption model.
@@
Furthermore, the Nikuradse dataset provided no evidence of test set leakage. This is a use case we foresee when scientists are trying to solve a mystery about their data while having a great deal of context to include potentially. Our experiments included <<100 words of context, but this context could include experimental details, instrument specifications, and related literature, pushing the limits of the LLM’s context window, or going beyond the context window in a retrieval-augmented generation scheme [58].
@@
The recent study by Shojaee and coworkers [89] demonstrated strong performance in generating scientific expressions using LLMs, using an approach that generated more structured outputs than our approach. Rather than rediscovering scientific laws as standalone expressions, they tasked the LLMs to “autocomplete” a structured Python function within the context of an optimization problem. They also designed new benchmark scenarios and datasets to highlight the strengths of this approach. This approach effectively leverages the strengths of language models in reasoning about well-defined problem statements. However, the extent to which its success can be attributed to prior exposure to similar questions during training remains unclear without a scratchpad to gain insights into its reasoning steps. A true blind test with these language models would be to rediscover an unpublished or recently published scientific law, using an LLM with a knowledge cutoff predating the seminal publication.
@@
Biasing the search space in a standard SR program can be challenging and requires advanced software and coding skills. Interdisciplinary work demands significant time and resources from researchers. We demonstrate how natural language interfaces in LLMs can help reduce some of these barriers by making program execution more accessible, even without expertise in software development. Well-crafted prompts empower language models to perform diverse tasks, allowing them to adapt to different contexts and objectives. We can guide these models through prompts to generate responses that align with our specific needs, whether solving complex problems, developing creative content, or analyzing data.
There are two ways to implement in-context learning in LLMs through prompts. One is the Few-Shot prompting method, where we condition the model on a few specific examples related to the task that help the model understand and perform the task more accurately. The other is the Zero-Shot prompting method (used here), where the output relies solely on template-based prompts without specific task examples, allowing the model to infer how to handle the task from general instructions. Our approach does not provide examples of symbolic regression procedures within the prompts. Instead, we guided the model to engage in free-form chain-of-thought reasoning about the context as it prepares suggestions for new equations.
@@ 5.1 System Design
Figure 20: Illustration of GPT-4 attempting symbolic regression. GPT-4 predicts expressions with optimized coefficients when passed a dataset for nitrogen adsorption on mica [60]. The Python code snippet from Prompt 2 output has been truncated to keep the figure concise. Note that the actual parameter values produced by running the code differ from what GPT-4 generates.
@@
5.2 Additional results
The following three datasets are from [25], carefully selected for evaluating SR algorithms for scientific data. The original data (Bode’s Law [10], Hubble’s Law [47], and Kepler’s Law [52]) were manually curated by [25] which we have used without modification.
Figure 23: Pareto fronts for Bode’s Law.
Figure 24: Pareto fronts for Hubble’s Law.
Figure 25: Pareto fronts for Kepler’s Third Law.